# Breaking Down the Barriers: Investigating Non-Expert User Experiences in Robotic Teleoperation in UK and Japan

### v1

@@ IV Experimental Material and Methodology
Participants were required to complete multiple questionnaires throughout the experiment. Initially, they completed a brief demographics questionnaire, which included age, gender, and experience with virtual reality, robots, and wearable gaming devices (e.g., Wiimote). These questions utilized a 5-point Likert scale to assess each participant’s perception and understanding. After teleoperating a robot, users answered two additional questionnaires. The first was the Single Ease Question (SEQ) [47], regarded as an effective end-of-task metric. The second was a raw NASA-TLX, described in Section II, employing a 7-point Likert scale instead of the 21-point scale, as research has shown that the 21-point scale does not enhance questionnaire reliability [48, 49].
@@
To explore the impact of the capabilities and limitations of the robotic hardware used for teleoperation, as well as the impact of demographics and previous experience, we deployed TELESIM on three different robots: the Rethink Robotics Baxter robot, the UR3 and the UR55e. Baxter and UR3 were used to conduct our experiments at the University of Glasgow in the UK, while, the Universal Robot 5e, at the Advanced Institute of Science and Technology (AIST) in Tokyo, Japan. This section explores the difference in setup and environment from the user survey developed in the UK. Due to regulations of AIST we were required to recruit our participants through an agency. Thus, we had access to a more significant reach than in the UK, reaching participants who were not involved in academia (i.e. undergraduate and postgraduate students and academics). Additionally, participants were financially compensated for participating in the user survey, unlike in the UK, where they participated for no monetary rewards. However, to keep our analysis consistent, we recruited participants within the same age range as described below:
@@ IV-A Robotic Setups in the UK and Japan
To facilitate an accurate comparison of the three robots employed in this study, a comprehensive analysis of their respective strengths and weaknesses is necessary. Table I provides a summary of the robots’ capabilities.
Baxter exhibits the longest reach at 1210 mm, followed by the UR5e with 850 mm, and the UR3 with 350 mm. However, Baxter’s extended reach introduces end effector position variability [52], a characteristic not observed in the Universal Robots. This variability is denoted as ”no stability” in Table I.
The primary distinctions between the two Universal robots lie in the UR5e’s extended reach, gripper configuration and control methodology. The UR3 was equipped with a modified Yale T42 gripper [45], measuring 21 cm in length when fully closed. Due to its size relative to the robot’s body, restrictive joint limits were implemented to prevent gripper-body collisions, consequently limiting the robot’s operational area. Conversely, the UR5e utilized a Robotiq 2F-140 gripper, measuring 23.3 cm when closed. For consistency, identical joint limits were applied to the UR5e; however, its longer reach mitigated the impact of these restrictions.
A notable difference between the grippers is the height differential between their open and closed states as mentioned before in Section III-C. The Robotiq gripper exhibits a 2.35 cm difference, while the T42 demonstrates has an 11 cm difference. This disparity increased the difficulty for users to estimate the appropriate gripper height for cube grasping with the UR3, as explained in Section IV-A.
The T42 gripper’s design introduced additional grasping challenges. As reported in Telesim [5], the Yale gripper’s limited closing force occasionally resulted in cube slippage during transit. This limitation was not observed with the Robotiq gripper. Lastly, the control methods differed between the robots. The UR3 was controlled via a Senseglove with an attached Vive tracker, while both the UR5e and Baxter utilized a Valve VR controller.
@@
For our Japan experiments, the TELESIM framework underwent minimal modifications, primarily involving the URDF file, which now loads a UR5e robot and replaces the T42 Gripper with a Robotiq 2F-140 gripper[46]. This change was required given the significant design flaws of the T42 gripper, as discussed in Section III-C. The Digital Twin was updated to reflect the new location and incorporate revised virtual safety settings, which constrain the motion planner from approaching certain areas due to safety considerations. Owing to the compatibility between Universal Robots 3 and 5e, we were able to utilize the same ROS2 controller developed during the initial TELESIM implementation.
Consistent with our previous UR3 experiment, the gripper’s collision parameters were configured based on its closed state to prevent hardware damage. To maintain analytical consistency, the joint position and speed limits applied to the UR5e were identical to those used in the UR3 experiment, despite the UR5e’s extended reach. Due to the extended reach, our previous cube placement was no longer suitable as the robot was able to reach all the location on our available table. Nevertheless, the experimental design aimed to preserve task difficulty by positioning cubes in a manner that required participants to adjust their body position to reach all cubes. This typically involved a single step in various directions, depending on the target cube. The cube arrangement and the step required to access the leftmost cube are depicted in Figure 5.
Figure 5: Overview of our experimental setup in AIST. The steam VR headset (1), as seen behind the robot and the user, acts as the world’s origin. The UR5e robot (2) in front is controlled by a Steam Index VR controller (3), in the hand of the user. There are 3 cubes setup in their starting position (4) in a triangular pattern similar to [5]. The empty square in the middle of the table (5), represent the location where the user should stack all of the cubes resulting in a tower of 3 cubes.
@@ V-A Experimental Results
Figure 6: Population percentage for each towers for Baxter in blue, the UR3 in orange and the UR5e in green
@@
Figure 6 depicts a steep performance improvement for the UR5e up to the 18th tower, followed by a plateau for the remaining seven towers. This plateau is attributed to the fact that only two participants managed to construct 21 and 25 towers, respectively. The exceptional performance of these participants suggests prior experience with robots or robotic teleoperation. However, the current questionnaire design, being self-reported, may not accurately capture this experience due to potential over- or under-estimation of skills by participants.
@@
To investigate the relationship between teleoperation performance and robot capabilities, additional statistics were collected, as shown in Figure 8. It is important to note that the teleoperation user surveys for Baxter and UR3 were conducted in the UK, while those for the UR5e were performed in Japan. The differences in experimental setup are detailed in Section IV.
@@
The Still in Place rate, representing the frequency of cubes remaining in their correct position without collapse, revealed a statistically significant advantage for the UR5e over Baxter (P <<< 0.01). This allows for the rejection of the null hypothesis that users achieved similar tower stability with Baxter and the UR5e. This conclusion is further supported by the difference in collapse rates, with the UR5e showing a statistically significant difference compared to Baxter (P <<< 0.1).
@@
Interestingly, the results of the Single Ease Question (SEQ) were statistically indistinguishable for Baxter and the UR5e (P >>> 0.01), despite the disparity in tower construction performance. Baxter achieved a mean SEQ score of 3.32 (SD = 1.27), the UR3 scored 2.19 (SD = 1.14), and the UR5e scored 3.72 (SD = 1.54). Higher scores indicate greater perceived accessibility. We hypothesise that the UR5e’s enhanced stability may have led users to perceive a higher tower construction potential than Baxter. Notably, participants were not informed of the maximum achievable tower count prior to completing the questionnaire. It appears that users estimated the theoretical maximum based on their experience, resulting in comparable SEQ scores for both robots despite performance differences. As expected, there is a clear relationship between the number of towers and a higher SEQ score, as can be seen in Figure 10
Figure 10: Regression plot showing relationship between SEQ and the amount of towers for each individual robots.
Our analysis primarily focused on the robots’ capabilities. However, it is important to note that the UR5e experiment was conducted in Japan, and we hypothesise that demographic factors may have influenced the results. The decision to conduct this user survey in Japan was based on research by Nam et al. [53], which suggests that user workload varies depending on the level of trust in a system, and Bartneck et al. [6], which indicates that participants from different countries exhibit varying levels of trust towards robots.
Japanese participants self-reported higher levels of experience with robots (mean 3.86 ± 1.27) compared to UK participants (mean 2.76 ± 1.24). This difference is statistically significant (P <<< 0.001) and can be attributed to the greater exposure of Japanese individuals to robots in their daily lives. Interestingly, Japanese participants reported less experience with wearable technology (mean 2.81 ± 1.47) compared to British participants (3.86 ± 0.87), a difference that is also statistically significant (P <<< 0.01). However, virtual reality (VR) experience was statistically similar between Japanese (mean 3.25 ± 1.38) and British (mean 2.97 ± 1.20) participants. These findings suggest that previous experience with robots and wearables varies depending on the user’s background, while VR experience remains consistent across both demographics, possibly due to the recent widespread adoption of VR technology [54].
It is important to note that we are unable to directly correlate previous experience with robots or wearables to teleoperation performance due to the differences in the systems operated by users. However, we were able to exclude previous VR experience as a contributing factor to performance differences.
@@ V-B Operator Workload
  • Physical: This subscale assesses the user’s physical workload during task execution. The UR3 exhibits a statistically significant difference compared to the other two robots, attributable to its distinct control method. The UR3 requires users to maintain a flat, extended hand position, unlike the VR controller, which allows for a more relaxed hand posture closer to the body.
@@
  • Effort: This subscale investigates the perceived effort spent by the user to complete the task. The UR3 shows the highest effort expenditure, which is statistically significant t the 95% level compared to Baxter and the UR5e. The consistently high effort scores across all robots corroborate the assertion that robot teleoperation is challenging for non-experts, as noted in previous research [55, 56].
  • Frustration: This subscale explores the user’s frustration and stress levels during the experiment. The UR3 leads with a statistically significant difference over Baxter and the UR5e, likely due to the disadvantages outlined in Section IV, particularly its limited range and weak gripper. Notably, this is the only subscale where participants utilized the full range of possible answers for all robots, suggesting that frustration tolerance is highly individual and may not be an ideal metric for evaluating teleoperation system performance.
@@ V-C Trust Towards Robots
    I would hate the idea that robots or artificial intelligence were making judgements about things.
Figure 12 reveals that Japanese users demonstrated higher trust in robots for questions NARS Q.1, NARS Q.2, and NARS Q.5. However, this difference is statistically significant only for NARS Q.5 (P<<<0.05). This general trend results in a statistically significant difference in total scores between the two demographics (P<<<0.05). This disparity may be attributed to the longer exposure of Japanese participants to robots in their daily lives, as noted by Bartneck et al. [6]. Additionally, Japanese culture’s lesser distinction between natural and artificial entities, influenced by Buddhist beliefs that do not discriminate against spirits in machines (unlike Christianity), may play a role [57, 6]. Media portrayal of robots may also contribute, with Western media often depicting robots as antagonists, while Japanese manga presents a more nuanced view where robots can assist in combating human-originated evil [6]. The significant difference in NARS Q.5 particularly supports this notion, as the concept of robots making decisions often leads to their rebellion in Western media [58].
Interestingly, our results appear to contradict the conclusion drawn by Bartneck et al. [6] that Japanese participants are less trusting than British participants. We hypothesize that this discrepancy may be due to the advancement and proliferation of robots and AI in daily life since the publication of this research work.
Finally, it is noteworthy that this difference in trust does not appear to affect teleoperation performance, as Japanese participants achieved the highest number of towers built. This outcome is primarily attributed to hardware differences, given that previous experience with VR-related technology is similar between Japanese and British participants, as explained in Section V-A.
@@ VI Conclusion and Future Work
This paper presents an extensive exploration of teleoperation’s impact on users through three user surveys, encompassing over 70 participants across two countries, utilizing three different robots and two control methods. The study employed NASA-TLX questionnaires to evaluate workload, NARS questionnaires to assess trust towards robots, and SEQ to gauge task difficulty. Additionally, the research investigated trust towards robots across different demographics.
@@
Experimental results demonstrate that participants achieved the highest number of towers built while using the UR5e. This superior performance is attributed to the UR5e’s enhanced stability and reach, allowing for better cube placement and reduced tower collapse during continued teleoperation. The UR5e also induced the least cognitive stress, while the combination of Senseglove and UR3 resulted in the highest physical strain and user frustration (cf. Figure 11).
@@
The findings in this paper corroborate previous research [55, 56] indicating that teleoperation is challenging for non-experts, characterized by high frustration and physical stress across all robot types. As anticipated, Japanese participants reported more VR experience than British participants. Conversely, Japanese participants reported less experience with wearable technology. The study demonstrated that prior VR experience does not significantly impact teleoperation performance, regardless of the operator’s background.
@@
Interestingly, Japanese participants exhibited higher trust towards robots compared to British participants, contradicting the conclusions drawn by Bartneck et al. [6]. These user surveys also validated the TELESIM framework by enabling international-scale teleoperation across multiple robotic platforms. In addition, the consistent results across demographics proved the user-friendliness of the TELESIM framework, along with its plug-and-play capability, as shown by its deployment on multiple robotic interfaces and end effectors.
@@
While this paper provides an in-depth analysis of teleoperation’s impact on users, it focuses solely on direct teleoperation with visual feedback. Other teleoperation methods mentioned in the literature, such as VR headset usage or telepresence robot control [16], were not explored. In future work, we aim to use these systems and investigate their impact on teleoperation success and user experience, following the methodology presented in this paper. However, our work on direct teleoperation is not completed. While our experiments seem to indicate that demographics do not play a part in the performance of teleoperation systems, we would need a wider breadth of demographics to show this conclusively. Furthermore, the many differences in hardware during our experiments make it difficult to accurately pinpoint the factors that impact the most, either the teleoperation performance or the user’s workload. Finally, the exact skillset required to perform teleoperation with minimal training beforehand remains an unsolved problem. The study successfully eliminated prior virtual reality experience as a confounding variable. However, a participant from Japan, who possesses expertise in motion retargeting for robotic teleoperation, demonstrated superior performance by completing the highest number of towers. This observation suggests that spatial abilities or other complex skills may influence task performance.

### v2

@@ IV Experimental Material and Methodology
Participants were required to complete multiple questionnaires throughout the experiment. Initially, they completed a brief demographics questionnaire, which included age, gender, and experience with virtual reality, robots, and wearable gaming devices (e.g., Wiimote). These questions utilized a 5-point Likert scale to assess each participant’s perception and understanding. After teleoperating a robot, users answered two additional questionnaires. The first was the Single Ease Question (SEQ)  [47], regarded as an effective end-of-task metric. The second was a raw NASA-TLX, described in Section II, employing a 7-point Likert scale instead of the 21-point scale, as research has shown that the 21-point scale does not enhance questionnaire reliability  [48, 49]. This research was validated by the University of Glasgow Ethics Committee (Application Number 300220026)
@@
To explore the impact of the capabilities and limitations of the robotic hardware used for teleoperation, as well as the impact of demographics and previous experience, we deployed TELESIM on three different robots: the Rethink Robotics Baxter robot, the UR3 and the UR5e. Baxter and UR3 were used to conduct our experiments at the University of Glasgow in the UK. At the same time, the Universal Robot 5e was at the National Institute of Advanced Industrial Science and Technology (AIST) Waterfront Center in Tokyo, Japan. This section explores the difference in setup and environment from the user survey developed in the UK. Due to time constraints, we were required to recruit our participants through an agency. Thus, we had access to a more significant reach than in the UK, reaching participants who were not involved in academia (i.e. undergraduate and postgraduate students and academics). Additionally, participants were financially compensated for participating in the user survey, unlike in the UK, where they participated for no monetary rewards. In addition, this research was internally approved by the Ergonomics Experiment Committee of the Life Sciences Experiment Management Office in the Environmental Safety Department of AIST with the following application number: 2023-1384. To keep our analysis consistent, we recruited participants within the same age range as described below:
@@ IV-A Robotic Setups in the UK and Japan
A comprehensive analysis of their strengths and weaknesses is necessary to help us better understand the three robots employed in this study. Table I summarizes the robots’ capabilities.
Baxter exhibits the longest reach at 1210 mm, followed by the UR5e with 850 mm and the UR3 with 350 mm. However, Baxter’s extended reach introduces end effector position variability  [52], a characteristic not observed in the Universal Robots. This variability is denoted as ”no stability” in Table I.
The primary distinctions between the two Universal robots lie in the UR5e’s extended reach, gripper configuration, and control methodology. The UR3 was equipped with a modified Yale T42 gripper  [45], measuring 21 cm when fully closed. Due to its size relative to the robot’s body, restrictive joint limits were implemented to prevent gripper-body collisions, consequently limiting the robot’s operational area. Conversely, the UR5e utilized a Robotiq 2F-140 gripper, measuring 23.3 cm when closed. For consistency, identical joint limits were applied to the UR5e; however, its longer reach mitigated the impact of these restrictions.
A notable difference between the grippers is the height differential between their open and closed states, as mentioned in Section III-C. The Robotiq gripper exhibits a 2.35 cm difference, while the T42 demonstrates an 11 cm difference. This disparity increased users’ difficulty in estimating the appropriate gripper height for cube grasping with the UR3, as explained in Section IV-A.
The T42 gripper’s design introduced additional grasping challenges. As reported in TELESIM  [5], the Yale gripper’s limited closing force occasionally resulted in cube slippage during transit. This limitation was not observed with the Robotiq gripper. Lastly, the robots’ control methods differed. The UR3 was controlled via a Senseglove with an attached Vive tracker, while both the UR5e and Baxter utilized a Valve VR controller.
@@
For our Japan experiments, the TELESIM framework underwent minimal modifications, primarily involving the URDF file, which now loads a UR5e robot and replaces the T42 Gripper with a Robotiq 2F-140 gripper [46]. This change was required given the significant design flaws of the T42 gripper, as discussed in Section III-C. The Digital Twin was updated to reflect the new location and incorporate revised virtual safety settings, constraining the motion planner from approaching certain areas due to safety considerations. Owing to the compatibility between Universal Robots 3 and 5e, we were able to utilize the same ROS2 controller developed during the initial TELESIM implementation.
Consistent with our previous UR3 experiment, the gripper’s collision parameters were configured based on its closed state to prevent hardware damage. To maintain analytical consistency, the joint position and speed limits applied to the UR5e were identical to those used in the UR3 experiment despite the UR5e’s extended reach. Due to the extended reach, our previous cube placement was no longer suitable, as the robot was able to reach all the locations on our available table. Nevertheless, the experimental design aimed to preserve task difficulty by positioning cubes in a manner that required participants to adjust their body position to reach all cubes. Depending on the target cube, this typically involves a single step in various directions. The cube arrangement and the step required to access the leftmost cube are depicted in Figure 5.
Figure 5: Overview of our experimental setup in AIST. The Steam VR headset (1) acts as the world’s origin, as seen behind the robot and the user. The UR5e robot (2) in front is controlled by a Steam Index VR controller (3). There are 3 cubes set up in their starting position (4) in a triangular pattern similar to  [5]. The empty square in the middle of the table (5) represents the location where the user should stack all the cubes, resulting in a tower of 3 cubes.
@@ V-A Experimental Results
Figure 6: Population percentage for each tower for Baxter in blue, the UR3 in orange, and the UR5e in green
@@
Figure 6 depicts a steep performance improvement for the UR5e up to the 18th tower, followed by a plateau for the remaining seven towers. This plateau is attributed to the fact that only two participants managed to build 21 and 25 towers, respectively. The exceptional performance of these participants suggests prior experience with robots or robotic teleoperation. However, the current questionnaire design, being self-reported, may not accurately capture this experience due to potential over- or under-estimation of skills by participants.
@@
Additional statistics were collected to investigate the relationship between teleoperation performance and robot capabilities, as shown in Figure 8. It is important to note that the teleoperation user surveys for Baxter and UR3 were conducted in the UK, while those for the UR5e were performed in Japan. The differences in experimental setup are detailed in Section IV.
@@
The Still in Place rate, representing the frequency of cubes remaining in their correct position without collapse, revealed a statistically significant advantage for the UR5e over Baxter (P <<< 0.01). This allows rejecting the null hypothesis that users achieved similar tower stability with Baxter and the UR5e. This conclusion is further supported by the difference in collapse rates, with the UR5e showing a statistically significant difference compared to Baxter (P <<< 0.1).
@@
Interestingly, the results of the Single Ease Question (SEQ) were statistically indistinguishable for Baxter and the UR5e (P >>> 0.01) despite the disparity in tower construction performance. Baxter achieved a mean SEQ score of 3.32 (SD = 1.27), the UR3 scored 2.19 (SD = 1.14), and the UR5e scored 3.72 (SD = 1.54). Higher scores indicate greater perceived accessibility. We hypothesize that the UR5e’s enhanced stability may have led users to perceive a higher tower construction potential than Baxter. Notably, participants were not informed of the maximum achievable tower count before completing the questionnaire. It appears that users estimated the theoretical maximum based on their experience, resulting in comparable SEQ scores for both robots despite performance differences. As expected, there is a clear relationship between the number of towers and a higher SEQ score, as can be seen in Figure 10
Figure 10: Regression plot showing the relationship between SEQ and the number of towers for each individual robot.
Our analysis primarily focused on the robots’ capabilities. However, it is important to note that the UR5e experiment was conducted in Japan, and we hypothesize that demographic factors may have influenced the results. The decision to conduct this user survey in Japan was based on research by Nam et al.  [53], which suggests that user workload varies depending on the level of trust in a system. Bartneck et al.  [6] indicate that participants from different countries exhibit varying levels of trust toward robots.
The Japanese participants self-reported higher levels of experience with robots (mean 3.86 ± 1.27) compared to the British participants (mean 2.76 ± 1.24). This difference is statistically significant (P <<< 0.001) and can be attributed to the greater exposure of the Japanese participants to robots in their daily lives. Interestingly, the Japanese participants reported less experience with wearable technology (mean 2.81 ± 1.47) compared to the British participants (3.86 ± 0.87), a difference that is also statistically significant (P <<< 0.01). However, virtual reality (VR) experience was statistically similar between the Japanese participants (mean 3.25 ± 1.38) and the British participants (mean 2.97 ± 1.20) participants. These findings suggest that previous experience with robots and wearables varies depending on the user’s background. In contrast, VR experience remains consistent across both demographics, possibly due to the recent widespread adoption of VR technology  [54].
It is important to note that we are unable to directly correlate previous experience with robots or wearables to teleoperation performance due to the differences in the systems users operate. However, we were able to exclude previous VR experience as a contributing factor to performance differences.
@@ V-B Operator Workload
  • Physical: This subscale assesses the user’s physical workload during task execution. The UR3 exhibits a statistically significant difference compared to the other two robots, attributable to its distinct control method. The UR3, controlled by the Senseglove, requires users to maintain a flat, extended hand position, unlike the VR controller, which allows for a more relaxed hand posture closer to the body.
@@
  • Effort: This subscale investigates the perceived effort spent by the user to complete the task. The UR3 shows the highest effort expenditure, which is statistically significant at 95% compared to Baxter and the UR5e. The consistently high effort scores across all robots corroborate the assertion that robot teleoperation is challenging for non-experts, as noted in previous research  [55, 56].
  • Frustration: This subscale explores the user’s frustration and stress levels during the experiment. The UR3 leads with a statistically significant difference over Baxter and the UR5e, likely due to the disadvantages outlined in Section IV, particularly its limited range and weak gripper. Notably, this is the only subscale where participants utilized the full range of possible answers for all robots. This suggests that frustration tolerance is highly individual and may not be an ideal metric for evaluating teleoperation system performance.
@@ V-C Trust Towards Robots
    I would hate the idea that robots or artificial intelligence were making judgments about things.
Figure 12 reveals that the Japanese users demonstrated higher trust in robots for questions NARS Q.1, NARS Q.2, and NARS Q.5. However, this difference is statistically significant only for NARS Q.5 (P<<<0.05). This general trend results in a statistically significant difference in total scores between the two demographics (P<<<0.05). This disparity may be attributed to the longer exposure of the Japanese participants to robots in their daily lives, as noted by Bartneck et al.  [6]. Additionally, Japanese culture’s lesser distinction between natural and artificial entities, influenced by Buddhist beliefs that do not discriminate against spirits in machines (unlike Christianity), may play a role  [57, 6]. Media portrayal of robots may also contribute, with Western media often depicting robots as antagonists. At the same time, Japanese manga presents a more nuanced view where robots can assist in combating human-originated evil  [6]. The significant difference in NARS Q.5 particularly supports this notion, as the concept of robots making decisions often leads to their rebellion in Western media  [58].
Interestingly, our results appear to contradict the conclusion drawn by Bartneck et al.  [6] that the Japanese participants are less trusting than the British participants. We hypothesize that this discrepancy may be due to the advancement and proliferation of robots and AI in daily life since the publication of this research work.
Finally, it is noteworthy that this difference in trust does not appear to affect teleoperation performance, as the Japanese participants achieved the highest number of towers built. This outcome is primarily attributed to hardware differences, given that previous experience with VR-related technology is similar between the Japanese and the British participants, as explained in Section V-A.
@@ VI Conclusion and Future Work
This paper presents an extensive exploration of teleoperation’s impact on users through three user surveys, encompassing over 70 participants across two countries, utilizing three different robots and two control methods. The study employed NASA-TLX questionnaires to evaluate workload, NARS questionnaires to assess trust toward robots, and SEQ to gauge task difficulty. Additionally, the research investigated trust towards robots across different demographics.
@@
Experimental results demonstrate that participants achieved the highest number of towers built using the UR5e. This superior performance is attributed to the UR5e’s enhanced stability and reach, allowing for better cube placement and reduced tower collapse during continued teleoperation. The UR5e also induced the least cognitive stress, while combining the Senseglove and UR3 resulted in the highest physical strain and user frustration (cf. Figure 11).
@@
The findings in this paper corroborate previous research  [55, 56] indicating that teleoperation is challenging for non-experts, characterized by high frustration and physical stress across all robot types. As anticipated, the Japanese participants reported more VR experience than the British participants. Conversely, the Japanese participants reported less experience with wearable technology. The study demonstrated that prior VR experience does not significantly impact teleoperation performance, regardless of the operator’s background.
@@
Interestingly, the Japanese participants exhibited higher trust towards robots compared to the British participants, contradicting the conclusions drawn by Bartneck et al.  [6]. These user surveys also validated the TELESIM framework by enabling international-scale teleoperation across multiple robotic platforms. In addition, the consistent results across demographics proved the user-friendliness of the TELESIM framework, along with its plug-and-play capability, as shown by its deployment on multiple robotic interfaces and end effectors.
While this paper provides an in-depth analysis of teleoperation’s impact on users, it focuses solely on direct teleoperation with visual feedback. Other teleoperation methods mentioned in the literature, such as VR headset usage or telepresence robot control  [16], were not explored. Moreover, additional comparison with the state of the art is difficult due to the disparity in the teleoperation task and control methods.
In future work, we aim to use these systems and investigate their impact on teleoperation success and user experience, following the methodology presented in this paper. However, our work on direct teleoperation is not completed. While our experiments seem to indicate that demographics do not play a part in the performance of teleoperation systems, we would need a wider breadth of demographics to show this conclusively. Furthermore, the many differences in hardware during our experiments make it difficult to accurately pinpoint the factors that impact the most, either the teleoperation performance or the user’s workload. A deeper investigation of the Senseglove and T42 gripper’s performance on the UR3 would enable a more direct comparison with the UR5e and further substantiate our hypothesis that hardware correlates with enhanced performance. This analysis would help elucidate whether the UR3’s suboptimal performance stems from gripper limitations, control methodology, or its restricted operational range.
@@
Finally, the exact skill set required to perform teleoperation with minimal training beforehand remains an unsolved problem. The study successfully eliminated prior virtual reality experience as a confounding variable. However, a participant from Japan, who possesses expertise in motion retargeting for robotic teleoperation, demonstrated superior performance by completing the highest number of towers. This observation suggests that spatial abilities or other complex skills may influence task performance.